OpenStudy (anonymous):

MAJOR HELP! verify the identity cotx/1+csc x = cscx -1/cotx

OpenStudy (mathstudent55):

is this it? \( \dfrac{\cot x}{1 + \csc x} = \dfrac{\csc x - 1}{\cot x} \)

OpenStudy (anonymous):

yes it is it :)

OpenStudy (anonymous):

\[\frac{ \cot x }{1+\csc x }*\frac{ \csc x-1 }{\csc x-1 }=\frac{ \cot x \left( \csc x-1 \right) }{ \csc ^{2}x-1 }\] \[use \csc ^{2} x-\cot ^{2}x=1 \] and get the solution

OpenStudy (anonymous):

how do i get the solution?

OpenStudy (anonymous):

its not multiplication they are equal to each other

OpenStudy (anonymous):

\[\csc ^{2}x-1=\cot ^{2}x\]

OpenStudy (anonymous):

\[=\frac{ \cot x \left( \csc x-1 \right) }{ \cot ^{2}x }=\frac{ \csc x-1 }{ \cot x }=R.H.S\]

OpenStudy (anonymous):

Thanks so much :)

OpenStudy (anonymous):

I have started with L.H.S and proved it equal to R.H.S

OpenStudy (mathstudent55):

Here's a different approach. By working on both sides, you can show the two sides to be equal. \( \dfrac{\cot x}{1 + \csc x} = \dfrac{\csc x - 1}{\cot x}\) \( \dfrac{ \dfrac{\cos x}{\sin x}}{1 + \dfrac{1}{\sin x}} = \dfrac{ \dfrac{1}{\sin x} - 1}{\dfrac{\cos x}{\sin x}} \) \( \dfrac{\sin x}{\sin x} \dfrac{ \dfrac{\cos x}{\sin x}}{1 + \dfrac{1}{\sin x}} = \dfrac{ \dfrac{1}{\sin x} - 1}{\dfrac{\cos x}{\sin x}} \dfrac{\sin x}{\sin x} \) \( \dfrac{ \cos x}{\sin x + 1} = \dfrac{ 1 -\sin x}{\cos x} \) \( (\cos x)(\sin x + 1)\dfrac{ \cos x}{\sin x + 1} = \dfrac{ 1 -\sin x}{\cos x} (\cos x)(1 +\sin x) \) \(\cos^2 x = 1 - \sin^2 x\) \( \cos^2 x = \cos^2 x\)